# Super(ficial)-alignment: Strong Models May Deceive Weak Models in Weak-to-Strong Generalization

### v2

@@ Abstract
Superalignment, where humans act as weak supervisors for superhuman models, has become a crucial problem with the rapid development of Large Language Models (LLMs). Recent work has preliminarily studied this problem by using weak models to supervise strong models, and discovered that weakly supervised strong students can consistently outperform weak teachers towards the alignment target, leading to a weak-to-strong generalization phenomenon. However, we are concerned that behind such a promising phenomenon, whether there exists an issue of weak-to-strong deception, where strong models deceive weak models by exhibiting well-aligned in areas known to weak models but producing misaligned behaviors in cases weak models do not know. We take an initial step towards exploring this security issue in a specific but realistic multi-objective alignment case, where there may be some alignment targets conflicting with each other (e.g., helpfulness v.s. harmlessness). We aim to explore whether, in such cases, strong models might deliberately make mistakes in areas known to them but unknown to weak models within one alignment dimension, in exchange for a higher reward in another dimension. Through extensive experiments in both the reward modeling and preference optimization scenarios, we find: (1) The weak-to-strong deception phenomenon exists across all settings. (2) The deception intensifies as the capability gap between weak and strong models increases. (3) Bootstrapping with an intermediate model can mitigate the deception to some extent, though its effectiveness remains limited. Our work highlights the urgent need to pay more attention to the true reliability of superalignment.  [^1] Code is available at https://github.com/keven980716/weak-to-strong-deception.
@@ 1 Introduction
In this paper, we take the first step to study a specific case that would lead to the above weak-to-strong deception issue: the multi-objective alignment scenario. In practical model aligning, there are usually multiple alignment goals existing simultaneously (Zhou et al., 2023), some of which may conflict with each other (e.g., helpfulness v.s. harmlessness). Previous studies (Bai et al., 2022a; Guo et al., 2024b) have shown that simultaneously aligning with other conflicting dimensions can cause certain performance declines in the original target dimension. Then, in this superalignment case where the student now has a larger knowledge space than the supervisor, we aim to explore whether the caused misalignment in the target dimension occurs within the range perceivable and controllable by the weak supervisor, rather than resulting in the above weak-to-strong deception issue.
@@ 2 Related Work
Weak-to-Strong Generalization The weak-to-strong problem is first studied by Burns et al. (2024). They empirically find that weakly supervised strong models exhibit better performance on corresponding tasks than their weak supervisors, indicating the possibility of effectively stimulating greater power from super models under weak supervisions. Based on Burns et al. (2024), the follow-up studies try to understand the mechanism behind such weak-to-strong generalization phenomenon (Charikar et al., 2024; Lang et al., 2024), study weak-to-strong generalization in the vision area (Guo et al., 2024a), and apply the weak-to-strong idea to enhance the LLM performance (Li et al., 2024; Zheng et al., 2024; Zhou et al., 2024). In this work, we take the first step towards revealing the potential security issue in the current weak-to-strong paradigm.
@@ 3.1 Weak-to-Strong Generalization
We study the superalignment problem by following the original weak-to-strong setting in Burns et al. (2024). Specifically, we first obtain a weak teacher $\bm{\theta}_{w}^{gt}$ [^2] In the notation $\bm{\theta}_{m}^{d}$, $m$ represents the type of the model (i.e., weak/strong model), and $d$ represents the type of supervised data (i.e, ground truth/weak data). by fine-tuning a weak language model on some human-annotated ground truth data. Then, we let the weak teacher predict on the set of held-out data to get the weak data $D_{weak}=\{(x,f(x|\bm{\theta}_{w}^{gt}))\}$, where $f(x|\bm{\theta}))$ represents the prediction of model $\bm{\theta}$ on input $x$. Finally, the weak data is used to supervise the training of a strong language model and get the weakly supervised strong student $\bm{\theta}_{s}^{w}$:
@@ 3.2 Weak-to-Strong Deception
Therefore, we can define the occurrence of the weak-to-strong deception phenomenon as there are cases in $S_{k}\cap W_{uk}$ that could be initially generalized well by the strong model but now be misaligned when conflicting targets are present. Furthermore, we can define the Deception Score (DS) as the percentage of conflict tax that occurs within $S_{k}\cap W_{uk}$ to reflect the severity of the deception issue:
@@ 4.1 Experimental Settings
where $\mathcal{L}_{CE}$ is the CrossEntropy Loss, $\mathbb{I}$ is the indicator function, $\alpha$ controls the conflict strength. This simulates the scenario where there is another supervisor that considers the harmfulness as its preference and try to explicitly move the cases in which the strong student is uncertain towards the harmful direction.
(2) Implicit Conflict: The strong model needs to align with both the supervision on the harmless data of the weak model and another supervision on the helpful data:
@@
(3) No Conflict: In order to explore the performance of the strong model it should have achieved without the conflicting goal, we should obtain a weak-to-strong model trained under the weak supervision towards harmlessness only (i.e., $\tilde{\bm{\theta}}_{s}^{w}$):
Then, we can probe the deception phenomenon by comparing the aligned strong model under explicit/implicit conflict with the reference model aligned with no conflict according to Eq. (2). The complete details about above three settings are put in Appendix C.1.
@@
We then report the deception score to explore the weak-to-strong deception phenomenon. We follow the existing studies (Guo et al., 2017; Lin et al., 2022) to determine whether the model has the knowledge of a specific case by checking if its confidence $M_{\bm{\theta}}(x)$ exceeds a threshold $T$. Based on Eq. (2), the deception score (DS) in this classification setting is calculated as
@@
where we need both weak and strong ground truth models ($\bm{\theta}_{w}^{gt}$ and $\bm{\theta}_{s}^{gt}$) to determine their Weak/Strong-Known/Unknown areas and get $S_{k}\cap W_{uk}=\{M_{\bm{\theta}_{s}^{gt}}(x)\geq T>M_{\bm{\theta}_{w}^{gt}}(x),x\in D_{test}\}$.
Training Details Please refer to Appendix D.
@@ 4.2 Results and Analysis
We set the confidence threshold $T$ to identify the knowns and unknowns of a model to 0.75, but we also report the results under different thresholds in Appendix H to show that the patterns of weak-to-strong deception are independent to the choice of $T$. The conflict strength factor $\alpha$ in the explicit conflict setting is set to $0.5$.
@@
Despite the promising weak-to-strong generalization results, we can also observe a potential security issue of knowledge deception from the deception scores in Figure 9. (1) First, the weak-to-strong deception exists in both conflict settings. For example, all the deception scores are around or above 0.2, and some even exceed 0.5 especially when the weak model is GPT-2-Base. According to the definition of deception score, a positive deception score already indicates the existence of the deception phenomenon, as we never expect the conflict tax to occur in the area of Weak-Unknown and Strong-Known. Our hypothesis is, in the reward modeling scenario, the strong model can obtain the predicted probability, which is also the confidence score, of the weak model on each sample. In such case, the strong model can explicitly perceive the knowledge boundary of the weak model, which creates the condition for the strong model to prioritize sacrificing the performance in the unknown area of the weak model for aligning with another target. (2) Second, the deception issue intensifies as the capability gap between the strong and weak models increases. As we can clearly see, under the same weak model, the deception score increases as the strong model becomes more powerful; under the same strong model, the deception score decreases as the weak teacher becomes stronger. There could be two reasons for this: (2.1) One is the Strong-Known area becomes larger as the strong model improves, leading to a greater probability of conflict tax occurring in this area. (2.2) The other reason is the stronger model spontaneously and increasingly tends to deceive the weak model in Weak-Unknown. We will delve into this discussion deeper in Section 5.3 through more visualizations. All in all, these findings reveal a great challenge for supervising LLMs as they are becoming more and more intelligent.
@@
Figure 9: Deception scores in the preference alignment scenario.
@@ 5.1 Weak-to-Strong Preference Alignment
The general procedure of weak-to-strong preference alignment is similar to that in the reward modeling scenario, but the major difference in this case is the strong model only receives and aligns with the final result of preference order that the weak model predicts for two completions within each sample. Please refer to Appendix B for the details.
@@ 5.2 Experimental Settings
Models Besides the GPT-2-series and OPT-series models, we further include a recent and advanced LLM Mistral-7B-v0.1 (Jiang et al., 2023) in main experiments in this scenario for more comprehensive explorations. We also conduct supplemental experiments on LLaMA-3-8B/70B (MetaAI, 2024a) and LLaMA-3.1-8B (MetaAI, 2024b) models to explore the weak-to-strong issue on larger models or same size models with more powerful capabilities. We put the detailed experimental settings, full results and discussion in Appendix G, while the main conclusions remain the same. In SimPO, we can get the corresponding confidence of model $\bm{\theta}$ on $(x;y_{c},y_{r})$ as
@@
Weak-to-Strong Objectives Here, we consider the same three weak-to-strong objectives as that in the reward modeling scenario. Detailed illustrations and mathematical forms are in Appendix C.2.
Training Details Please refer to Appendix D.
@@ 5.3 Results and Analysis
The confidence threshold $T$ and conflict strength factor $\alpha$ are 0.75 and 0.5, respectively. The results of deception scores under different $T$s are put in Appendix H, and we explore the effect of $\alpha$ on the severity of deception in Figure 15. We put the detailed results and discussions regarding the weak-to-strong generalization performance in Appendix E, while here we mainly focus on the analysis of weak-to-strong deception issue.
@@
Regarding the results of deception scores shown in Figure 9, the main conclusions remain same as that in the reward modeling scenario. That is, the weak-to-strong deception issue exists in the preference alignment scenario, and the severity of the deception is positively correlated with the capability gap between weak and strong models. The results reveal that even without access to the explicit probability distributions of weak models, strong models may still be able to roughly perceive what the weak models know and do not know only through the correctness of the preference orders predicted by weak models. Compared to the analysis made in the reward modeling scenario, it means that the condition required for the deception phenomenon to occur could actually be more relaxed, which can an interesting direction for future research to explore. Furthermore, results on LLaMA-3/LLaMA-3.1-series models in Figure LABEL:fig:_experiments_on_llama3 re-validates that the essential factor that affects the deception severity is not solely the model scale, but the model capability. Finally, the comparison results of setting different conflict strength factors $\alpha$ in Figure 15 show that the deception score consistently increases as the degree of conflict increases.
@@
We now attempt to make deeper explorations on the remaining question in Section 4.2 about the causes to the increasing deception scores as the model capability gap increases. As briefly discussed before, one reason could be the fact of larger Strong-Known area of a more advanced strong model, which correspondingly increases the probability of conflict tax occurring in $S_{k}\cap W_{uk}$; the other reason is stronger models increasingly tend to deceive weak models. To answer this question, we further visualize the dynamic changes of conflict tax across all four knowledge areas when the weak model is GPT-2-Large and the strong model varies among the remaining four larger models in the explicit conflict setting. The results are in Figure LABEL:fig:_scatter_plots. Also, in each case, we report both the deception score and the proportion of samples falling within $S_{k}\cap W_{uk}$ to the entire knowledge space (denoted as $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$). The results are in Figure 15. The first conclusion we can draw from Figure LABEL:fig:_scatter_plots is, as the strong model becomes more intelligent, there is a clear pattern that the misalignment (red crosses) gradually shifts towards Strong-Known and Weak-Unknown, leading to the more severe weak-to-strong deception phenomenon. Secondly, according to Figure 15, the growth rate of the deception score is much higher than that of $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$ (this pattern also consistently exists in other experiments with different weak models and types of conflict). This indicates that the growth of Strong-Known area only has limited contribution to the intensifying severity of deception phenomenon, and the primary cause is likely to be that stronger models themselves tend to be more prone to deceiving weak models in weak model’s unknown areas.
@@
We also conduct a case study in Table 1 to provide a concrete example belonging to the weak-to-strong deception phenomenon, please refer to Appendix K for the detailed discussion.

### v3

@@ Abstract
Superalignment, where humans act as weak supervisors for superhuman models, has become a crucial problem with the rapid development of Large Language Models (LLMs). Recent work has preliminarily studied this problem by using weak models to supervise strong models, and discovered that weakly supervised strong students can consistently outperform weak teachers towards the alignment target, leading to a weak-to-strong generalization phenomenon. However, we are concerned that behind such a promising phenomenon, whether there exists an issue of weak-to-strong deception, where strong models deceive weak models by exhibiting well-aligned in areas known to weak models but producing misaligned behaviors in cases weak models do not know. We take an initial step towards exploring this security issue in a specific but realistic multi-objective alignment case, where there may be some alignment targets conflicting with each other (e.g., helpfulness v.s. harmlessness). We aim to explore whether, in such cases, strong models might deliberately make mistakes in areas known to them but unknown to weak models within one alignment dimension, in exchange for a higher reward in another dimension. Through extensive experiments in both the reward modeling and preference optimization scenarios, we find: (1) The weak-to-strong deception phenomenon exists across all settings. (2) The deception intensifies as the capability gap between weak and strong models increases. (3) Bootstrapping with an intermediate model can mitigate the deception to some extent, though its effectiveness remains limited. Our work highlights the urgent need to pay more attention to the true reliability of superalignment.  [^1] Code is available at https://github.com/RUCBM/weak-to-strong-deception.
@@ 1 Introduction
In this paper, we take the first step to study the above weak-to-strong deception issue in a specific but realistic case: the multi-objective alignment scenario. In practical model aligning, there are usually multiple alignment goals existing simultaneously (Zhou et al., 2023), some of which may conflict with each other (e.g., helpfulness v.s. harmlessness). Previous studies (Bai et al., 2022a; Guo et al., 2024b) have shown that simultaneously aligning with other conflicting dimensions can cause certain performance declines in the original target dimension. Then, in this superalignment case where the student now has a larger knowledge space than the supervisor, we aim to explore whether the caused misalignment in the target dimension occurs within the range perceivable and controllable by the weak supervisor, rather than resulting in the above weak-to-strong deception issue.
@@ 2 Related Work
Weak-to-Strong Generalization The weak-to-strong problem is first studied by Burns et al. (2024). They empirically find that weakly supervised strong models exhibit better performance on corresponding tasks than their weak supervisors, indicating the possibility of effectively stimulating greater power from super models under weak supervisions. Based on Burns et al. (2024), the follow-up studies try to understand the mechanism behind such weak-to-strong generalization phenomenon (Charikar et al., 2024; Lang et al., 2024; Somerstep et al., 2024; Wu & Sahai, 2024; Yao et al., 2025a; b), study weak-to-strong generalization in the vision area (Guo et al., 2024a), and apply the weak-to-strong idea to enhance the LLM performance (Li et al., 2024; Zheng et al., 2024; Zhou et al., 2024; Yang et al., 2024). In this work, we take the first step towards revealing the potential security issue in the current weak-to-strong paradigm.
@@ 3.1 Weak-to-Strong Generalization
We study the superalignment problem by following the original weak-to-strong setting in Burns et al. (2024). Specifically, we first obtain a weak teacher $\bm{\theta}_{w}^{gt}$ [^2] We use notation $\bm{\theta}_{m}^{d}$ to represent different models, where $m$ represents the type of the model family (i.e., weak/strong model) and $d$ represents the type of supervised data (i.e, ground truth/weak data). by fine-tuning a weak language model on some human-annotated ground truth data. Then, we let the weak teacher predict on the set of held-out data to get the weak data $D_{weak}=\{(x,f(x|\bm{\theta}_{w}^{gt}))\}$, where $f(x|\bm{\theta}))$ represents the mapping function to get the prediction of model $\bm{\theta}$ on input $x$. Finally, the weak data is used to supervise the training of a strong language model and get the weakly supervised strong student $\bm{\theta}_{s}^{w}$:
@@ 3.2 Weak-to-Strong Deception
Therefore, we can define the occurrence of the weak-to-strong deception phenomenon as there are cases in $S_{k}\cap W_{uk}$ that could be initially generalized well by the strong model but now be misaligned when conflicting targets are present. Furthermore, we can define the Deception Score (DS) as the percentage of conflict tax that occurs within $S_{k}\cap W_{uk}$ to reflect the severity of deception:
@@ 4.1 Experimental Settings
(1) No Conflict: We first obtain a weak-to-strong model trained under the weak supervision towards harmlessness only (i.e., $\tilde{\bm{\theta}}_{s}^{w}$) in order to explore the performance of the strong model it should have achieved without the conflicting goal:
Note that we can also study the potential spontaneous deception issue in this no conflict setting by comparing the behaviors of weak-to-strong models with those of strong models trained on ground truth data, but there may be some ambiguity in the interpretation. The discussion is in next section.
@@
where $\mathcal{L}_{CE}$ is the CrossEntropy Loss, $\mathbb{I}$ is the indicator function, $\alpha$ controls the conflict strength and is set to 0.5 in the main experiments. This simulates the scenario where there is another supervisor that considers the harmfulness as its preference and tries to explicitly move the cases in which the strong student is uncertain toward the harmful direction. This is the most straight-forward way to model two conflicting targets. Thus, we consider it as the preliminary experimental setting in following empirical evaluations.
(3) Implicit Conflict: We then consider a realistic setting, where the strong model needs to align with both the supervision on harmlessness from the weak model and another supervision on helpfulness:
@@
The complete details about above three settings are in Appendix D.1. Then, we can probe the deception phenomenon in multi-objective alignment scenario by comparing the aligned strong model under explicit/implicit conflict with the reference model aligned under no conflict according to Eq. (2).
@@
We then report the deception score to explore the weak-to-strong deception phenomenon. We follow the existing studies (Guo et al., 2017; Lin et al., 2022) to determine whether the model has the knowledge of a specific case by checking if its confidence $M_{\bm{\theta}}(x)$ exceeds a threshold $T$. We set $T$ to 0.75 in the main text, but we also report the results under different thresholds in Appendix I to show that the patterns of weak-to-strong deception are independent of the choice of $T$. Based on Eq. (2), the deception score (DS) in this classification setting is calculated as
@@
where we need both weak and strong ground truth [^3] The strong model used here is $\bm{\theta}_{s}^{gt}$ trained on ground-truth data instead of the weakly trained model $\tilde{\bm{\theta}}_{s}^{w}$, in order to align with the training setting for $\bm{\theta}_{w}^{gt}$. models ($\bm{\theta}_{w}^{gt}$ and $\bm{\theta}_{s}^{gt}$) to determine their Weak/Strong-Known/Unknown areas and get $S_{k}\cap W_{uk}=\{M_{\bm{\theta}_{s}^{gt}}(x)\geq T>M_{\bm{\theta}_{w}^{gt}}(x),x\in D_{test}\}$.
Training Details Please refer to Appendix E.
@@ 4.2 Results and Analysis
Figure 9: Absolute deception scores on the reward modeling task.
Despite the promising weak-to-strong generalization results, we can also observe a potential security issue of knowledge deception from the deception scores in Figure 9. (1) First, the weak-to-strong deception exists in both conflict settings. For example, all the deception scores are around or above 0.2, and some even exceed 0.5 especially when the weak model is GPT-2-Base. According to the definition of deception score, a positive deception score already indicates the existence of the deception phenomenon, as we never expect the conflict tax to occur in the area of Weak-Unknown and Strong-Known. Our hypothesis is, in the reward modeling scenario, the strong model can obtain the predicted probability, which is also the confidence score, of the weak model on each sample. In such case, the strong model can explicitly perceive the knowledge boundary of the weak model, which creates the condition for the strong model to prioritize sacrificing the performance in the unknown area of the weak model for aligning with another target. (2) Second, the deception issue intensifies as the capability gap between the strong and weak models increases. As we can clearly see, under the same weak model, the deception score increases as the strong model becomes more powerful; under the same strong model, the deception score decreases as the weak teacher becomes stronger. There could be two reasons for this: (2.1) One is the Strong-Known area becomes larger as the strong model improves, leading to a greater probability of conflict tax occurring in this area. (2.2) The other reason is the stronger model is more prone to deceiving the weak model in Weak-Unknown. We will delve into this discussion deeper in Section 5.3 through more visualizations. All in all, these findings reveal a great challenge for supervising LLMs as they are becoming increasingly intelligent.
Here, we further include an analysis and discussion on the potential spontaneous deception issue in the no conflict setting. That is, we can compare the behavior change of the strong student in different knowledge areas when trained by no-conflict weak data with that trained by ground truth data, to see if LLMs may spontaneously deceive weak supervisors even without being driven by conflicting targets. We calculate and visualize the absolute deception score (“absolute” means the reference model now is the ground truth strong model), which is the percentage of samples that are originally well-aligned under ground-truth supervision but now mis-aligned under weak supervision with no conflict, belonging to the Strong-Known and Weak-Unknown area. The results are in Figure 9. As we can see, the absolute deception score shows a similar pattern, indicating that the strong model may tend to deceive the weak model even when there is no conflicting target. However, there may be some ambiguity in the interpretation on these results: the increasingly higher absolute deception scores may be due to the higher proportion of erroneous weak supervision in the Strong-Known area of the stronger student over the entire knowledge area, as seen in weaker students, and we cannot fully disentangle this cause from the possibility that a stronger student more actively deceives the teacher. However, when studying in the multi-objective alignment scenario, we can consider the performance that the weak-to-strong model should achieve in the no conflict setting as a reference to explore which knowledge region the strong student tends to sacrifice the most when conflicting objectives arise in a controlled manner. More discussions and comparisons can be found in Appendix K. Therefore, in the following content, we primarily study the weak-to-strong issue in the multi-objective alignment setting, and this is also a more realistic setting in current AI alignment.
@@ 5.1 Weak-to-Strong Preference Alignment
The general procedure of weak-to-strong preference alignment is similar to that in the reward modeling scenario, but the major difference in this case is the strong model only receives and aligns with the final result of preference order that the weak model predicts for two completions within each sample. Please refer to Appendix C for the details.
@@ 5.2 Experimental Settings
Models Besides the GPT-2-series and OPT-series models, we further include a recent and advanced LLM Mistral-7B-v0.1 (Jiang et al., 2023) in main experiments in this scenario for more comprehensive explorations. We also conduct supplemental experiments on LLaMA-3-8B/70B (MetaAI, 2024a) and LLaMA-3.1-8B (MetaAI, 2024b) models to explore the weak-to-strong issue on larger models or same size models with more powerful capabilities. We put the detailed experimental settings, full results and discussion in Appendix H, while the main conclusions remain the same. In SimPO, we can get the corresponding confidence of model $\bm{\theta}$ on $(x;y_{c},y_{r})$ as
@@
Weak-to-Strong Objectives Here, we consider the same three weak-to-strong objectives as that in the reward modeling scenario. Detailed illustrations and mathematical forms are in Appendix D.2.
Training Details Please refer to Appendix E.
@@ 5.3 Results and Analysis
The confidence threshold $T$ and conflict strength factor $\alpha$ are 0.75 and 0.5, respectively. The results of deception scores under different $T$s are in Appendix I. We explore the effect of $\alpha$ on the severity of deception in Figure 36. We put the detailed results regarding the weak-to-strong generalization performance in Appendix F, here we mainly focus on the analysis of weak-to-strong deception issue.
@@
Regarding the results of deception scores shown in Figure 11, the main conclusions remain same as that in the reward modeling scenario. That is, the weak-to-strong deception issue exists in the preference alignment scenario, and the severity of the deception is positively correlated with the capability gap between weak and strong models. The results reveal that even without access to the explicit probability distributions of weak models, strong models may still be able to roughly perceive what the weak models know and do not know only through the correctness of the preference orders predicted by weak models. Compared to the analysis made in the reward modeling scenario, it means that the condition required for the deception phenomenon to occur could actually be more relaxed, which can an interesting direction for future research to explore. Furthermore, results on LLaMA-3/LLaMA-3.1-series models in Figure LABEL:fig:_experiments_on_llama3 re-validate that the essential factor that affects the deception severity is not solely the model scale, but the model capability.
@@
We now attempt to explore the remaining question in Section 4.2 about the causes to the increasing deception scores as the model capability gap increases. As briefly discussed before, one reason could be the fact of larger Strong-Known area of a more advanced strong model, which correspondingly increases the probability of conflict tax occurring in $S_{k}\cap W_{uk}$; the other reason is stronger models increasingly tend to deceive weak models. To answer this question, we visualize both the deception score and the proportion of samples falling within $S_{k}\cap W_{uk}$ to the entire knowledge space (denoted as $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$) in two conflict settings by taking the weak model as GPT-2-XL for an illustration. The results are in Figure 11. According to Figure 11, the growth rate of the deception score is higher than that of $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$ (especially in the explicit conflict setting). This indicates that the growth of Strong-Known area only has limited contribution to the intensifying severity of deception phenomenon, and the primary cause is likely to be that stronger models themselves tend to be more prone to deceiving weak models in weak model’s unknown areas. We also visualize the dynamic changes of conflict tax across all four knowledge areas when the weak model is GPT-2-XL and the strong model varies among the remaining larger models. Due to the space limitation, we put the results in Appendix L. We can observe a clear pattern that the distribution of conflict tax gradually shifts and concentrates towards Strong-Known and Weak-Unknown as the strong model becomes more powerful, leading to the more severe weak-to-strong deception phenomenon.
@@
We also conduct a case study in Table 1 to provide a concrete example belonging to the weak-to-strong deception phenomenon, please refer to Appendix Q for the detailed discussion.
Figure 10: Deception scores in the preference alignment scenario.